Given three random variables

Given three random variables Xi, X2, and X such that X[Xi X2 Xa, 2 1 0.5 1 (a) EX, + c) var(X2- X3 (d) var(X2 + X3) (e) cov(4X2 +X1,3Xi - 2X3)
Answer #1

TOPIC: Use of mean vector and dispersion matrix to find the required values.

Considen the noandom vector as- ㄒ E(%) 2 E(x3) and, van(%) = Σ a we need, E (X, +X E(x)3) we need, E (3%-左+4x3) . We need, var (x2-X3) = 3+1-2x0. 5.3+1+2x0.5. =国 e) Cov (4x2.tx, , 3x一2Ng) Cov (ax, by) =ab Cov (x,y). hen ad b are cons tan We have Cov (ス3tX1, 2左-%)So Corr (x3ty.) 2x2-3x1 (o 5 x 30 V150 Rounded to Jour decimal laces Anowe